DI Question 2 4.6 pt The interpretation of the coefficient on In(Pi) in the model In(Q)-9.76-0.5 In(Pi) + ui, where Qi is the quantity demanded for a good and Pi is the price of the good, is: a 1% increase in P is associated with a 50% decrease in Q O an increase in P by one unit is associated with a 0.5 units decrease in Q. 0 a 1% increase in P is associated with a 0.5% decrease in Q. 0 a 1% increase in P is associated with a 0.5 units decrease in Q.
Answer #1

Case 1: A 1% increase in P

lets say p=100% initially So lnQ=9.76-0=9.76 So Q=17326.63

then 1% increase(p=101%) So lnQ=9.76-0.004975=9.7550 So Q=17240.64

So %change = (17326.63-17240.64) / 17326.63 * 100 = 0.5% So option C

Case 2 if we consider an increase in p by one unit then there will be 0.5 units decrease in lnQ

(not in Q) So option B is wrong